### v1

# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers From Driving Video

Abstract
We introduce scenario-based cognitive status identification in older drivers from Naturalistic driving videos and large vision models. In recent times, cognitive decline, including Alzheimer’s disease (AD) and mild cognitive impairment (MCI), is often underdiagnosed due to the time-consuming and costly nature of current diagnostic methods. By analyzing real-world driving behavior captured through in-vehicle systems, this research aims to extract “digital fingerprints” that correlate with functional decline and clinical features of MCI and AD. Moreover, modern large vision models can draw meaningful insights from everyday driving patterns of older patients to early detect cognitive decline. We propose a framework that uses large vision models and naturalistic driving videos to analyze driver behavior, classify cognitive status and predict disease progression. We leverage the strong relationship between real-world driving behavior as an observation of the current cognitive status of the drivers where the vehicle can be utilized as a “diagnostic tool”. Our method identifies early warning signs of functional impairment, contributing to proactive intervention strategies. This work enhances early detection and supports the development of scalable, non-invasive monitoring systems to mitigate the growing societal and economic burden of cognitive decline in the aging population.
keywords:
Naturalistic driving, Large Vision model, Cognitive decline, driving scenario
[…]
[label1]organization=Department of Electrical & Computer Engineering,addressline=Iowa State University, city=Ames, state=IA, country=USA
[…]
[label4]organization=Department of Mechanical Engineering,addressline=Iowa State University, city=Ames, state=IA, country=USA
1 Introduction
Cognitive decline in the aging population is a critical health concern that manifests across a spectrum of stages. The pre-dementia stage, commonly referred to as mild cognitive impairment (MCI), is characterized by measurable cognitive impairment with minimal or no functional changes in daily living. Individuals at this stage, often in their 70s, may exhibit early signs of cognitive decline without severe disruptions to independence. As the condition progresses to dementia, particularly Alzheimer’s disease (AD), more severe cognitive and functional impairments emerge, ultimately diminishing the ability to live independently. While MCI is not exclusive to AD, it often serves as an early indicator of the disease and highlights the need for early detection and intervention.
This study focuses on individuals exhibiting MCI due to AD, as determined by clinical characteristics. The timely identification of such individuals provides an opportunity to slow disease progression, improve quality of life, and reduce the societal burden associated with advanced stages of dementia. Traditional methods for diagnosing cognitive decline are often time-consuming, costly, and reliant on clinical visits or complex neuroimaging techniques, which may not be practical for widespread use [1, 2]. To address this, our research explores an innovative approach using real-world driving behavior as a passive diagnostic tool.
Prior research on cognitive decline has explored a variety of methods to detect early changes in individuals at risk of AD. Neuroimaging techniques, such as magnetic resonance imaging (MRI) [3] and positron emission tomography (PET), have been widely used to identify structural and functional brain changes associated with MCI and AD. These methods are often paired with biomarkers, such as amyloid-beta and tau proteins, to improve diagnostic accuracy. However, the high cost and limited accessibility of these techniques hinder their widespread application.
Wearable sensor technologies [4, 5] have also gained traction in recent years. Studies have demonstrated the utility of devices that monitor gait patterns, heart rate variability, and sleep quality to assess functional and cognitive impairments. Machine learning models [6, 7] applied to sensor data have shown promise in predicting the progression of cognitive decline. However, these approaches often require participant compliance and are limited to specific contexts, such as controlled environments or clinical trials.
Traditional methods like road tests and driving simulators are commonly used to assess fitness-to-drive [8] but are limited to controlled conditions and face challenges related to objectivity, availability, and cost-effectiveness. To address these limitations, driving research has shifted towards naturalistic assessments using Sensor devices [9] installed in participants’ personal vehicles, enabling the collection of real-world driving behavior data.
Previous studies [10, 11] have demonstrated that everyday driving behavior, such as reduced driving distances, fewer unique destinations, and smaller driving spaces, is associated with symptomatic Alzheimer’s disease (AD). However, limited research has investigated the impact of preclinical AD on driving behavior [12], highlighting a critical gap in understanding early-stage cognitive decline.
In contrast, our approach capitalizes on passive monitoring of real-world driving behavior, which offers a non-invasive, cost-effective, and scalable solution. By combining driving video data with advanced computer vision techniques, we aim to bridge the gap between traditional clinical diagnostics and real-world applications, providing an innovative pathway for early detection of cognitive decline.
2 Methodology
2.1 RWRAD Dataset
This dataset was collected in collaboration with the University of Nebraska Medical Center and encompasses comprehensive neuropsychological assessments, clinical tests, and lab results to investigate cognitive impairment and its progression.
Table 1: Study participant characteristics
Variable Range or Counts Average Std. Dev.
Age (years) Range: 65 - 90 75.4 6.0
Sex Female: 75
Male: 61
Race White: 125
Black: 9
American Indian: 1
Other: 1
Income \$0 - \$100,000 \$43,026.43 \$27,855.35
Not indicated: 86
Employment Employed Full/Part Time: 19
Retired: 107
Volunteering: 9
Keeping house: 1
Driving Experience (years) Range: 45 - 78 58.7 6.6
Weekly Driving Range: 5 - 300 miles (8 - 483 km) 85.9 miles (138.2 km) 75.2 miles (121.0 km)
MoCA Score Range: 15 - 30 24.3 3.4
Participants were classified using established criteria for mild cognitive impairment (MCI), which identifies individuals with cognitive changes that do not yet produce major functional impairments. Table 1 shows the detailed demographic and cognitive layout of the participants. Key classification methods include:
  • 1.
    Jak criteria: [13] Two tests in one domain scoring more than 1 standard deviation (SD) below the mean, offering higher longitudinal accuracy but detecting fewer MCI cases.
  • 2.
    Peterson criteria: [14] One test in one domain scoring more than 1.5 SD below the mean, detecting more MCI cases but with reduced longitudinal accuracy.
MCI classifications primarily identify amnestic impairments, often associated with short-term and episodic memory, though impairments in other domains, such as language and visuospatial skills (atypical presentations), are increasingly recognized. As cognitive impairment progresses, it often becomes a multi-domain process, ultimately impacting functional abilities and transitioning to the dementia stage of clinical Alzheimer’s disease (AD). The study design for collecting in-lab data focused on assessing functional abilities and activities through driving behavior in various scenarios. Key questions included identifying participants’ characteristics and health status, determining cognitive impairments and their probable link to AD, evaluating co-morbid sources of cognitive decline, and assessing overall health, cognitive function, mobility, vision, and signs of dependence, depression or quality-of-life decline.
To address challenges associated with compliance and self-reporting accuracy in cognitively impaired participants, a study partner was recruited for each participant. Data collection included contributions from both participants and study partners, with comparative analyses used to ensure data reliability and robustness. This longitudinal dataset provides a valuable resource for studying the progression of cognitive decline and functional impairment.
Clinical evaluation of MCI due to AD is primarily a rule-out process aimed at determining whether the primary impairment is due to another type of dementia. However, the majority of people with AD ($\sim$60%) also have other types of dementia [15]. While this study does not exclude individuals with co-morbid non-AD dementias, clinical assessments are designed to capture these data. Participants recruited into this study are included because they exhibit primary AD. It is important to note that AD presents a complex and heterogeneous landscape, making its study both challenging and critical.
In addition to clinical diagnosis and functional assessments, composite cognitive scores can offer a quantitative summary of neuropsychological performance across multiple domains. A composite cognition score (COGSTAT) integrates standardized measures from several well-established neuropsychological tests. This approach captures impairments in memory, language, visuospatial skills, and executive function—core domains often affected in AD and MCI. The formula for COGSTAT is as follows:
$\text{COGSTAT}=350$ $+\frac{C-38.7}{11.1}+\frac{L-26.2}{3.5}+\frac{R-10.08}{3.2}$ (1)
$+\frac{B-4.4}{2.4}+\frac{R_{cs}-31}{3.7}+\frac{R_{rs}-15.7}{5.7}+\frac{T-46.1}{32.6}$
Where:
  • 1.
    $C$ = Controlled Oral Word Association test (continuous value)
  • 2.
    $L$ = Judgment of Line Orientation test
  • 3.
    $R$ = Auditory Verbal Learning test: 30-minute recall score
  • 4.
    $B$ = Number of errors in Benton Visual Retention test
  • 5.
    $R_{cs}$ = Rey-O Complex Figure test: copy score
  • 6.
    $R_{rs}$ = Rey-O Complex Figure test: 30-minute recall score
  • 7.
    $T$ = Trail Making test Part B: time in seconds
While such neuropsychological testing provides a comprehensive assessment of cognitive abilities, it often requires in-person clinical visits, participant compliance, and substantial time and resources — all of which may be challenging for individuals in the early stages of AD. In this context, analyzing naturalistic driving video offers a promising, non-invasive alternative to capture subtle behavioral cues of cognitive decline in real-world settings. Studies show that behavioral patterns in everyday driving can be an alternative and effective approach to detect early-stage AD, complementing traditional cognitive testing and potentially enabling earlier intervention.
2.2 Video Data Pre-pocessing
Figure 1: Omaha most transited route database GIS map (Omaha, NE)
Figure 2: Breakdown of missing and redundant video in RWRAD dataset
Figure 3: Cognitive score (a) COGSTAT, (b) MoCA distribution across subjects ($N=69$). We combined the AD and MCI subjects into a unified class“AD-aging” and performed two-class classification with the “Normal-aging”.
Figure 4: Preparation of video samples for specific scenarios. The top row represents the rollout of all driving videos and drives across subjects. These video files from different subject blocks are filtered based on scenarios and used as input to the Large Vision model.
Figure 5: Overview of the video-to-embedding generation. Given, input video segments $\mathcal{V}=\{V_{k}^{(D_{i},Sc_{j})}\}\mid i=1,\dots,d;\;j=1,\dots,k;\;k=1,\dots,n\}$ from different scenarios $Sc_{j}\in\mathcal{S}$, a vision model $F_{\theta}$ maps each $v_{i}$ to an embedding $Z_{i}=F_{\theta}(V_{i})\in\mathbb{R}^{m}$. The embedding space is analyzed to assess inter-group separation using a distance metric. Scenario $Sc_{1}$ exhibits greater class-wise separation than $Sc_{2}$, i.e., $\text{Dist}_{Sc_{1}}>\text{Dist}_{Sc_{2}}$, indicating higher discriminative potential for identifying cognitive states. Selected embeddings are passed to a classifier $H(Z_{i})\rightarrow\hat{Y}_{i}$ for downstream cognitive classification.
Data was collected during a 3-month baseline - including driving video, actigraphy, sleep log, driving behavior and in-lab data by the University of Nebraska Medical Center in Omaha. Ninety legally licensed and active, aged between 65 - 90 years old participants were recruited with a study partner near Omaha, NE. They were recruited through community outreach, including flyers, local news, and community engagement [16]. Over a 2-year period, driving data were collected in 3-month intervals using custom-installed “Black Boxes” in each participant’s personal vehicle. These systems passively recorded driving behavior throughout each trip, capturing synchronized forward roadway and cabin video alongside sensor data from GPS, speed, accelerometer, gyroscope, and throttle. All sensor data were processed at a 1Hz sampling rate. All participants provided informed consent in accordance with institutional protocols.
While analyzing the collected driving video data, we encountered some missing and corrupted video clips that required filtering before analysis. Fig. 2 presents the breakdown of the processed driving data for our analysis. While processing the in-vehicle Blackbox video data, we identified a total of 520,662.7 minutes of driving footage, segmented into pure drive duration (72.6%), blackframe drive duration (4.1%), and missing drive duration (23.3%). For downstream analysis, only valid segments—comprising both pure and black-frame video—were retained. Segments with missing or corrupted data were excluded to ensure input quality and consistency for driver behavior modeling and cognitive health assessment tasks.
Subjects included in this study were required to pass at least a predefined fraction of the Omaha most transited route road segments during the three-month baseline period. This criterion was applied to ensure adequate video data coverage for each participant and to make sure we have diverse driving scenario exposure in the dataset. By focusing on these criteria, our framework extracts meaningful insights into the cognitive status of aging drivers, correlating observed behaviors with functional decline. For our analysis, a total of $N=69$ subjects were utilized for video-based analysis using Year-1 data, after excluding the subjects without class labels and drivers outside the Omaha region. Figure 3 shows the distribution of COGSTAT scores across three aging groups— AD, MCI, and Normal aging. Among 69 subjects, 35 drivers (26 with MCI and 9 with Alzheimer’s Disease) were categorized as “AD-aging”, while 34 were classified as “Normal-aging”. This balanced distribution enabled a robust comparison between normal and AD aging groups in terms of cognitive test scores and driving behavior. Subjects with Alzheimer’s consistently show the lowest scores, while normal-aging individuals cluster at the higher end. MCI subjects span the intermediate range, overlapping with both groups. It also highlights the transitional nature of cognitive decline and the inherent classification challenge.
To ensure consistency and relevance in our analysis, we focused on short segments. We selected roadway segments where a drive can be $\leq 0.31$ mile or $\leq 500$ meters and defined it as a “route segment” (discussed in 3.1). These segments provide concentrated snapshots of driver behavior under controlled road conditions. This approach minimizes variability due to changes in driving context or prolonged periods of observation, which may dilute critical behavioral signals. Additionally, the time of day and traffic conditions (peak vs. off-peak hours) were considered to account for environmental stressors that could uniquely challenge drivers with Alzheimer’s clinical syndrome (ACS).
2.2.1 Scenario-specific Video Sample
The overall pipeline for scenario-specific video sample preparation is illustrated in Fig. 4. Let $\mathcal{D}={D_{1},D_{2},\dots,D_{d}}$ be the set of subjects that participated in this study. Each subject $D_{i}$ performs a set of drives {${dr_{1},dr_{2},\dots,dr_{n}}$} where $n$ is the maximum number of drives per subject. Each drive is further segmented into short video clips ${V_{1},V_{2},\dots,V_{m}}$, where $m$ is the maximum number of video clips per drive. These video clips are recorded under different driving scenarios $\mathcal{S}={Sc_{1},Sc_{2},\dots,Sc_{s}}$. From these clips, a subset of scenario-specific video samples $V_{k}^{(D_{i},Sc_{j})}$ is filtered for each subject-scenario pair $(D_{i},Sc_{j})$ where $k$ is the index of the sampled video clip. This resulted in a quality dataset of labeled video samples that were used as inputs to the large vision model in the next stage. Additionally, the hierarchical rollout of video samples across subjects, drives, and scenarios ensures a structured and scalable representation across varying cognitive profiles and driving contexts.
2.2.2 Large Vision Model-based Framework
We feed the video samples into a pre-trained Large Vision Transformer model [17] to perform the video-to-vector generation. The extracted vector representations are stored in Amazon AWS Cloud for further analysis. The goal of this framework is to transform raw driving videos into discriminative vector representations for downstream classification of cognitive aging groups. This is achieved via the following four-step pipeline: The framework follows four main steps: The entire framework includes four steps as follows:
Step 1: Video-to-embedding Generation
Each video sample $V_{k}^{(D_{i},Sc_{j})}$ is processed through a pre-trained large vision model $F_{\theta}$ to obtain a fixed-size vector space:
$Z_{k}=F_{\theta}(V_{k}^{(D_{i},Sc_{j})}),\quad Z_{k}\in\mathbb{R}^{m}$
where $Z_{k}$ represents the vector representation of each video sample $V_{k}^{(D_{i},Sc_{j})}$ in an $m$-dimensional latent space.
Step 2: Dimensionality Reduction
To interpret the structure of the vector space and the separability of the aging groups in each scenario setting, we apply a dimensionality reduction function $g:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ (e.g., PCA [18], t-SNE [19]) to obtain:
$\tilde{Z}_{k}=g(Z_{k}),\quad\tilde{Z}_{i}\in\mathbb{R}^{n},\text{ where }n\in\{50,100,200,\dots\}$
This step is performed in a way to preserve the majority of the information embedded in the high-dimensional video representation while mapping it into a lower-dimensional vector space. It allows us to observe how the visual features related to the driving behavior in different scenarios are clustered in a low-dimensional space and select the prominent scenarios for the next steps. As shown in Fig. 5, a demo reduced 2D $(n=2)$ representations are plotted to visualize group separability within each scenario.
Step 3: Discriminative Scenario Selection
To evaluate the relevance of a scenario $Sc_{j}\in\mathcal{S}$ for classification, we compute a pairwise distance metric (e.g., centroid, Euclidean) over the embeddings belonging to different classes (e.g., Normal, AD). The scenario with the highest inter-group distance is selected:
$Sc^{*}=\arg\max_{Sc_{j}\in\mathcal{S}}\ \text{Dist}_{\text{avg}}(Z_{\text{class}_{normal}}^{(Sc_{j})},Z_{\text{class}_{AD}}^{(Sc_{j})})$
where $Z_{\text{class}_{normal}}^{(Sc_{j})}$ and $Z_{\text{class}_{AD}}^{(Sc_{j})}$ denote the set of embeddings under scenario $Sc_{j}$ for class Normal and AD respectively.
Step 4: Classification using RF Model
The scenario-specific embeddings are then used to train a Random Forest (RF) [20] classifier $H:\mathbb{R}^{m}\rightarrow\mathcal{Y}$, where $\mathcal{Y}$ is the set of class labels:
$\hat{Y}_{k}=H(Z_{k})$
Final model performance is assessed using metrics such as $\mathbf{(a)}$, precision$\mathbf{(P)}$, recall$\mathbf{(R)}$ and $\mathbf{(F1)}$ under multiple sampling strategies (e.g., random vs. driver-level separation) to ensure generalizability. We report average classification accuracy under two sampling strategies:
  • 1.
    Random Sampling: A total of $n$ video samples are randomly selected from the full dataset for training and testing, without enforcing subject- or scenario-level separation. This strategy evaluates the model’s performance under unconstrained sampling conditions.
  • 2.
    Driver-level Separation (DLS): To evaluate the generalizability of the proposed framework across different individuals, we employed a leave-$k$-drivers-out cross-validation strategy [21], where $k$ drivers are excluded from the training set and reserved exclusively for testing. This procedure is repeated over $r$ independent runs, each with a different subset of left-out drivers. Finally, the results across these runs are averaged to get the performance metrics $\mathbf{(a,P,R,F1)}$. It ensures robustness and mitigates variance due to subject-specific biases.
This setup allows us to evaluate both intra-subject consistency and inter-subject generalization in the learned representations.
Algorithm 1 Scenario-Based Video Analysis
1:Input: Set of video samples $\mathcal{V}={V_{k}^{(D_{i},Sc_{j})}}$ for subject $D_{i}$ in scenario $Sc_{j}$
2:Output: Trained classifier $H$, scenario relevance ranking ${\text{Dist}_{\text{avg}}(Sc_{j})}$
3:for each scenario $Sc_{j}\in\mathcal{S}$ do
4:     Initialize $\mathcal{E}_{Sc_{j}}\leftarrow\emptyset$triangleright Store embeddings
5:     for each subject $D_{i}\in\mathcal{D}$ do
6:         for each video $V_{k}^{(D_{i},Sc_{j})}$ do
7:              $Z_{k}\leftarrow F_{\theta}(V_{k}^{(D_{i},Sc_{j})})$triangleright Video-to-vector
8:              $\mathcal{E}_{Sc_{j}}\leftarrow\mathcal{E}_{Sc_{j}}\cup{Z_{k}}$
9:         end for
10:     end for
11:     ~$\tilde{\mathcal{E}}_{Sc_{j}}\leftarrow g(\mathcal{E}_{Sc_{j}})$triangleright Dimensionality reduction, $g:\mathbb{R}^{m}\to\mathbb{R}^{n}$
12:     Compute $\text{Dist}_{\text{avg}}(Sc_{j})\leftarrow\text{Distance}(\tilde{\mathcal{E}}^{\text{normal}},\tilde{\mathcal{E}}^{\text{mci}})$
13:end for
14:$Sc^{*}\leftarrow\arg\max_{Sc_{j}}\ \text{Dist}_{\text{avg}}(Sc_{j})$triangleright Most discriminative scenario
15:Train RF classifier $H:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ using $\mathcal{E}_{Sc^{*}}$ and corresponding labels
16:return $H$, scenario ranking ${\text{Dist}_{\text{avg}}(Sc_{j})}$
3 Experimental Setup
3.1 Scenario-based Roadway Selection
To investigate the relationship between driving behavior and cognitive status, we selected two distinct types of roadway scenarios where older drivers commonly exhibit prominent visual or maneuver-based patterns. These patterns can serve as interpretable behavioral cues for identifying cognitive decline. The selected driving scenarios for our video-based analysis are:
  • 1.
    Freeway-interchange: merging or diverging behavior in ramp entries or exit segments. e.g., JFK N to I-80 E Ramp, I-480 W to Dodge W Ramp.
  • 2.
    Interstate: high-speed, lane-keeping scenarios on long uninterrupted roads. e.g., Omaha I-80 W, I-80 E.
(a) Interstate segments
(b) Freeway-interchange segments
Figure 6: GIS map - (a) Interstate and (b) Freeway-interchange scenarios. The relevant route segments were filtered out from the Omaha most transited route database shown in Figure 1.
We utilize Omaha most transited route database for filtering the roadway scenarios. From different federal functional categories (e.g., Local, Major collector, Minor collector etc.), we observe and select the most prominent scenario combination for this study. Figure 6 shows the GIS map highlighting the two roadway scenarios for our video analysis. We considered a single driving video having $\leq 0.31$ mile or $\leq 500$ meter length as a route segment. Omaha most transited route database consists of 24,106 route segments. In total, we filter out $964$ unique route segments, consisting of $632$ freeway-interchange segments (highlighted in green in Figure 6(b)) and $332$ interstate segments (highlighted in blue in Figure 6(a)). Table 2 summarizes the key statistics – a balance of the selected data and driving exposure of the cognitive groups. Additionally, we selected the route segments that maximize the number of drives for both cognitive groups. The avg. trips per unique subject column indicates that each driver has driven on the chosen roadway types at least 3.5 times, ensuring sufficient behavioral representation per subject.
To interpret the driving exposure data distribution, we define several key variables below:
  • 1.
    $N_{u}$: Number of unique trips by the Normal-aging subjects
  • 2.
    $M_{u}$: Number of unique trips by the subjects with MCI
  • 3.
    $A_{u}$: Number of unique trips by the subjects with AD
  • 4.
    $T_{u}$: Total unique trips across all subject groups
    $T_{u}=N_{u}+M_{u}+A_{u}$
    It does not include repeated trips which means multiple driving sessions recorded on the same drive segment by the same individuals
  • 5.
    $\overline{T}_{u}$: Total number of trips including repeated ones (non-unique).
    It includes repeated trips which means multiple driving sessions recorded on the same drive segment by the same individual.
  • 6.
    $T_{avg}$: Average number of trips per subject in a specific drive segment. It gives an approximation that drivers, on average, drove more on Interstate than on Freeway interchanges.
Table 2: Trip details by roadway types
Scenario Normal MCI AD Total Total Avg. trips per
categories trips($N_{u}$) trips($M_{u}$) trips($A_{u}$) trips($T_{u}$) trips($\overline{T}_{u}$) subject($T_{avg}$)
Fwy-int. 8136 10134 919 19189 68048 3.55
Interstate 7123 7894 780 15797 69620 4.41
$N_{u}$: unique trips by Normal group; $M_{u}$: unique trips by MCI group; $A_{u}$: unique trips by AD group; $T_{u}$: total unique trips; $\overline{T}_{u}$: total non-unique trips; Fwy-int.: Freeway interchange.
We leveraged “unique drives” vs. “repeated drives” per subject across all the route segments to indicate statistical reliability and diversity in driving exposure by the two cognitive groups. The average trips per subject suggest that each participant had substantial driving trips with the selected roadway scenario with a significant amount of video segments for our analysis.
Figure 7 shows a boxplot distribution of speed limits and the route distance for both scenarios. As we observe, the freeway-interchange segments tend to have lower speed limits and shorter distances, which works as a transitional driving scenario, involving challenging driving behavior and fast decision-making. In contrast, the interstate segments exhibit higher speed limits and longer travel distances, which indicate an uninterrupted highway driving scenario with relaxed driving behavior.
Figure 7: Boxplot distributions of the speed limit and route distance for Interstate and Freeway-interchange roadways. The middle line shows the median speed limit and median route distances.
  • 1.
    Median speed limits:
    Fwy-interchange scenario, $S_{\text{fwy}}=37.5$ mph;
    Interstate scenario, $S_{\text{int}}=60$ mph
  • 2.
    Median distances:
    Fwy-interchange scenario, $D_{\text{fwy}}=0.12$ miles;
    Interstate scenario, $D_{\text{int}}=0.20$ miles
These characteristics offer a strong contrast between the two scenarios, reflecting distinct driving demands and cognitive load associated with each scenario.
3.2 Executing Video-based Model
To evaluate our video-based embedding framework, we conducted experiments across two distinct driving scenarios that capture both emergency maneuvers and straight-path driving behaviors. The aim was to analyze whether these movement types could be effectively distinguished through video-derived embeddings.
The experiments were conducted under the following evaluation protocols:
  • 1.
    Random Sampling: Video segments are randomly split into training and testing sets without enforcing driver-level separation.
  • 2.
    Driver-level Separation: A leave-$k$-drivers-out cross-validation strategy is employed, where $k=5$ drivers were excluded during training and used exclusively for testing. This protocol was repeated over $r=3$ independent runs with different left-out driver groups, and the final metrics were averaged to capture subject variability.
In our experiments, the parameters are set as: (1) Input image size (960x752x3), (2) Embedded dimension (6144x1), (3) Frame Rate—$\{1,10\}$, (3) Low-dimension size— $n\in\{50,100,200\}$, and (4) Distance Metrics—L2 and centroid-distance. These experiments provide a comprehensive understanding of how video-derived embeddings behave under varying sampling conditions and data splits and lay the foundation for downstream classification tasks.
4 Results
4.1 Model Performance on Different Scenarios
Table 3: Model performance on two roadway types
Scenario Accuracy Precision Recall F1 score Sampling
categories $\mathbf{(a)}$ $\mathbf{(P)}$ $\mathbf{(R)}$ $\mathbf{(F1)}$ method
Fwy-int. 71.03% 0.7597 0.7728 0.7662 random
Interstate 55.12% 0.5404 0.5333 0.5368 random
$\Delta$ 15.91% 0.1293 0.1395 0.1344 random
Fwy-int. 69.81 ± 2.51% 0.7378 ± 0.0371 0.7624 ± 0.0671 0.7505 DLS
Interstate 52.17 ± 1.93% 0.5920 ± 0.0636 0.5997 ± 0.0771 0.5958 DLS
$\Delta_{\text{ds}}$ 17.64% $\uparrow$ 0.1458 $\uparrow$ 0.1627 $\uparrow$ 0.1547 $\uparrow$ DLS
$\Delta$: difference w/o DLS, $\Delta_{\text{ds}}$: difference w/ DLS, Fwy-int.: Freeway interchange, DLS: Driver-level Separation
Table 3 reveals the performance of the video-based framework trained to identify two cognitive aging-groups (Normal-aging vs. AD-aging) across two scenario categories—Freeway interchange and Interstate. The $\Delta$ row (without driver-level separation) indicates significant separation on all metrics $\mathbf{(a,P,R,F1)}$ and validates the selection of these two contrastive scenarios to classify aging groups. The performance difference (grey colored rows) between the two scenarios indicates that there are prominent vision features in these specific scenarios that can successfully distinguish the aging classes solely by looking at the driving video features.
Furthermore, the driver-level separation yields slightly widened separation, which can be observed on all metrics $\mathbf{(a,P,R,F1)}$ in the $\Delta_{\text{ds}}$ row. It surpassed the performance difference achieved in the without Driver-level Separation case; for example, the accuracy difference went up from $\Delta$ = 15.91% to $\Delta_{\text{ds}}$ = 17.64%, indicating a 1.73% increment in performance difference. The $\Delta_{\text{ds}}$ aligns with achieving a significant performance difference on two contrastive scenarios to classify aging groups while there is no “Subject-level overlapping”. It also validates that the Fwy-int. scenario includes more prominent driving movement features compared to its counterpart. On the other hand, the Interstate scenario performs significantly worse, yielding lower accuracy and F1-score. Therefore, this observation emphasizes the Freeway interchange as a more behaviorally informative and cognitively sensitive driving scenario, where complex maneuvers, short distances and lower speed variability have more nuanced driver responses.
We compared our results with a recent study [22] that used a multi-modal (demographics and sleep data) framework to detect cognitive decline in older drivers. Our method outperformed both without Driver-level Separation ([22] accuracy 68.64% vs. our Fwy-int. accuracy 71.03%) and with Driver-level Separation ([22] accuracy 70.48% vs. our Fwy-int. max accuracy 72.32%) cases on the RWRAD dataset. This highlights the effectiveness of scenario-specific vision features as non-invasive digital biomarkers for cognitive aging classification. Moreover, another recent study by Danda et al. [23] found that amyloid-positive older adults exhibited higher average speeds, intermittent speed variability and increased acceleration on freeway on-ramp driving (similar to Fwy-int. scenario) suggesting less adaptive driving. In clinical research, Amyloid is a protein that accumulates abnormally in the brain and forms plaques linked to the development of AD. Patients with cognitive impairment who test amyloid-positive are typically diagnosed as having AD, whereas amyloid-negative patients are generally excluded from an AD diagnosis [24, 25]. These findings confirm that Freeway ramp driving requires increased cognitive load involving behavioral reactions to rapidly appearing traffic and abrupt speed shifts, combining perception with motor control [23]. Additionally, they underscore our results of higher classification accuracy ($\mathbf{a}$ = 71.03%) and F1-score ($\mathbf{F1}$ = 0.6712) in Freeway interchange scenarios compared to the Interstate settings ($\mathbf{a}$ = 55.12%, $\mathbf{F1}$ = 0.5368) reported in Table 3. Although their approach combined driving patterns with physiological signals, it validates the relevance of using contrastive and challenging roadway scenarios in cognitive classification tasks. Therefore, it aligns with our findings that the Freeway interchange scenario provides more informative and strong behavioral cues that the vision model can leverage to better separate cognitive aging groups.
4.2 Percentage of Misclassification
Figure 8: Percentage of misclassification vs. - (a) COGSTAT score and (b) MoCA score. The relevant route segments were filtered out from the Omaha most transited route database shown in Figure 1.
In this section, we assess the feasibility of our modeling approach by calculating the percentage of misclassification for all $N=69$ subjects (including normal and AD). For each subject $k$, we denote $n_{\text{test}}^{(k)}$ represent the number of times they appeared in the test set across 10 resamples, and $n_{\text{error}}^{(k)}$ denote the number of times they were misclassified by the RF classifier. The subject-specific misclassification rate is then calculated as:
[…]
Figure 8 displays the relationship between the misclassification percentage and Cognitive scores (COGSTAT or MoCA). An overall trend observed in both plots is that subjects labeled as AD-aging (MCI and AD diagnoses) tend to exhibit higher misclassification percentages compared to the Normal-aging drivers. This is evident in the middle cognitive score ranges (COGSTAT: 580–650; MoCA: 12–20), where the misclassification is higher. It suggests that the middle zones correspond to a transition region where the classifier shows greater uncertainty, due to the overlapping feature space between early-stage impairment and normal aging. It underscores the anticipated challenge of reliably distinguishing subtle cognitive impairments/MCI from normal aging. On the other hand, subjects with low (COGSTAT $<550$ and MoCA $<8$) and high (COGSTAT $>700$ and MoCA $>20$) cognitive scores consistently yield lower misclassification rates. This behavior indicates that the classifier learned prominent feature representations for cognitively distinct individuals in these regions and it is able to classify the aging groups more confidently. Additionally, it reinforces the validity of using COGSTAT and MoCA as effective biomarkers in well-separated diagnostic states. Overall, these trends are consistent with the model’s performance as reported in Table 3. The agreement between global performance metrics and localized misclassification patterns supports the interpretation that the model’s predictive success is not random but rather driven by the prominent vision signals yielded from a certain cognitive state. Subjects with well-defined cognitive profiles— clear AD aging (Alzheimer/strong MCI) or normal aging are more reliably classified. However, uncertainty increases in transitional cognitive states, reflecting appropriate model sensitivity to the complexity of early-stage classification. This behavior supports the validity of our approach and suggests that further improvement may require incorporating auxiliary clinical signals to resolve the transition zone uncertainty. Overall, this subject-level metric captures individual classification stability and uncertainty at the subject level rather than aggregated classification accuracy alone.
4.3 Limitations and Future Works
The current setting evaluates only two types of roadways – freeway interchanges and interstates. Future work can include a broader set of driving scenarios (e.g., urban intersections, rural roads) to generalize findings across different traffic and environmental conditions. The driver-level separation was limited to 3-fold cross-validation with a fixed number of test subjects. Although this approach evaluates inter-subject generalization, the small fold count and limited test pool can lead to variance in performance estimation. Recruiting more diverse drivers and conducting longitudinal assessments would strengthen generalization. The use of in-vehicle video data raises potential concerns regarding privacy. Future implementations should address these with robust data management.
Another notable limitation is the relatively limited availability of high-quality in-vehicle video data per cognitive group. The model’s misclassification rates, particularly in the transitional region, are likely influenced by this data scarcity. We believe that acquiring a larger and higher-quality dataset with enriched video samples for both cognitively normal and AD groups will allow the model to learn more distinct and generalizable feature representations. This would be especially critical for improving classification accuracy in cases with confusing behavioral features.
5 Discussion
The results suggest that scenario-based analysis plays a significant role in video-based cognitive assessment. By segmenting driving video data into semantically distinct scenarios, our approach enables targeted analysis of driving behavior under specific cognitive demands. In the scope of this work, freeway interchange scenarios yield significantly higher classification performance compared to interstate scenarios under driver-level separation. Therefore, certain driving scenes inherently capture richer behavioral variance, which large vision models can leverage to distinguish between Normal vs. AD aging patterns. This approach offers a low-cost, scalable alternative to traditional cognitive screening methods, enabling passive and non-intrusive cognitive monitoring. It has direct implications for the healthcare industry, where early detection of cognitive decline remains a costly and time-consuming process. Moreover, the ability to extract digital biomarkers from routine driving behavior transforms personal vehicles into affordable diagnostic tools. This reduces patient burden while supporting continuous and longitudinal health assessment. The approach holds promise in providing an efficient digital health solution for aging populations in early detection of cognitive decline and screening in underserved or rural areas with limited access to healthcare facilities.
6 Acknowledgements
This work was supported by the National Institutes of Health (NIH), the National Institute on Aging (NIA 5R01AG17177-18), and the University of Nebraska Medical Center Mind & Brain Health Labs. The views expressed in this paper are those of the authors alone and not the NIH or NIA. We thank our research team for coordinating this project. Due to personally identifying information (PII), the data are not available online. The authors declare no potential conflict of interest.
7 Appendix